OpenStudy (katielong):

factorise 17x"+32x-20=0

jhonyy9 (jhonyy9):

what is the exponent of the first x ?

OpenStudy (katielong):

\[17x ^{2}\]

OpenStudy (katielong):

the original question is: solve these simultaneous equations \[4x+4=y\] \[x ^{2}+y ^{2}=36\] and so far i have got to: \[17x^{2}+32x-20\]

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[(x-x_1)(x-x_2)=0\]

OpenStudy (katielong):

?

OpenStudy (unklerhaukus):

apply the quadratic formula

OpenStudy (katielong):

ahhhh i see

OpenStudy (unklerhaukus):

\[17x^{2}+32x-20=0\]\[a=17\\b=32\\c=-20\]

OpenStudy (katielong):

yup im with you now, thanks

OpenStudy (katielong):

\[\frac{ -32+\sqrt{2384} }{ 34 }\] am i right so far?

OpenStudy (unklerhaukus):

well almost \[x_{1.2}=\frac{ -32\pm\sqrt{2384} }{ 34 }\](you need plus and minus, because there are two roots)

OpenStudy (katielong):

ah yup, thanks alot, got it now :)

OpenStudy (unklerhaukus):

so you can only simplify the sqrt a little bit

OpenStudy (unklerhaukus):

[the solutions are not Very simple ]

OpenStudy (unklerhaukus):

(but that is mainly because there are not many lines that intersect circles at 'nice' points )

OpenStudy (katielong):

yeah, the solutions are like -5.508967349....

OpenStudy (unklerhaukus):

i would leave √149 in , without approximating